# KVmix: Gradient-Based Layer Importance-Aware Mixed-Precision Quantization for KV Cache

### v1

@@ 5 Conclusion
This paper addresses the memory bottleneck issue posed by the KV Cache during the inference of LLMs by proposing a novel mixed quantization method, KVmix. KVmix achieves layer importance-aware mixed-precision quantization by leveraging KV weight gradients. It also provides dynamic long-context optimization to reduce memory usage while preserving generation quality. Experiments show that KVmix, with an average of 2.19 bits for Keys and 2.38 bits for Values, reduces inference memory usage by 4.9$\times$ and increases throughput by 5.3$\times$, while maintaining near-lossless inference quality. KVmix offers flexibility and ease of use, enabling the adjustment of quantization strategies based on specific application scenarios, thus effectively balancing model accuracy, inference efficiency, and memory usage. However, it has not yet achieved real-time mixed quantization during inference. Future work will explore lightweight mechanisms for real-time bit adjustments in mixed-precision quantization, enabling KVmix to dynamically respond to changing inference demands while maintaining efficiency.
Appendix A KV Weight Analysis
Fig. 7 presents the heatmaps of the KV weights ($W_{k}$ and $W_{v}$) for the Llama 3-8B-Instruct and Mistral-7B-Instruct-v0.3 models. For different layers of various models, the $W_{k}$ and $W_{v}$ exhibit significant differences. The differences are consistent with the conclusions in Section 2.2.
@@ Appendix A KV Weight Analysis
Figure 7: Projection matrix weights of K and V across different model layers for Llama 3-8b-Instruct (top) and Mistral-7B-Instruct-v0.3 (bottom). The heatmap visualizes the projection weights by sampling every sixth element in the matrices. Additionally, the weight values along the axes have been normalized. "Norm" represents the L2 norm of the weight matrix for each layer, while "Range" indicates the range of values within each layer’s weight matrix.
Appendix B KVmix Profiler Algorithm
In the Algorithm B, we provide the workflow of the KVmix profiler, including the process from data processing, KV importance analysis, and quantization bit configuration. When using the KVmix profiler, selecting 20 to 30 prompts is sufficient to yield reliable importance analysis results. It typically takes only 10 to 15 minutes to complete on a single NVIDIA RTX 4090 GPU.
@@ Appendix B KVmix Profiler Algorithm
Model $model$, inputs $input\_ids$, $attention\_mask$, $labels$, layers $num\_layers$, datasets $datasets$
Quantization bits $quant\_bits$
for dataset in datasets do
     prompts $\leftarrow$ randomSample(dataset, n);  $\triangleright$ Sample n prompts
     inputs $\leftarrow$ tokenize(prompts);  $\triangleright$ Prepare inputs
     model $\leftarrow$ loadFullModel();  $\triangleright$ Load full-precision model
@@
     kScores, vScores $\leftarrow$ calcImportance(model, inputs, num_layers);  $\triangleright$ Compute importance
@@
     kLayers, vLayers $\leftarrow$ classify(kScores, vScores, num_layers);  $\triangleright$ Classify layers
@@
     quantBits $\leftarrow$ setBits(kLayers, vLayers);  $\triangleright$ Set quantization bits
     del model;  $\triangleright$ Free memory
     model $\leftarrow$ loadQuantModel(quantBits);  $\triangleright$ Load quantized model
     preds $\leftarrow$ infer(model, dataset);  $\triangleright$ Inference
     save(preds);  $\triangleright$ Save results
     del model;  $\triangleright$ Free memory
end for
@@
Function calcImportance(model, inputs, num_layers):  $\triangleright$ Compute importance scores
kScoresAll, vScoresAll $\leftarrow$ [], [];
for input in inputs do
     kScores, vScores $\leftarrow$ [], [];
     for layer in range(num_layersdo
         loss $\leftarrow$ model.forward(input);  $\triangleright$ Compute loss
@@
         kGrad $\leftarrow$ grad(loss, kParams(layer));  $\triangleright$ Gradient for Key
@@
         vGrad $\leftarrow$ grad(loss, vParams(layer));  $\triangleright$ Gradient for Value
         kScores.append(kGrad.norm());  $\triangleright$ Append Key score
         vScores.append(vGrad.norm());  $\triangleright$ Append Value score
         clearMemory();  $\triangleright$ Free memory
     end for
     kScoresAll.append(kScores);
     vScoresAll.append(vScores);
end for
kScoresMean $\leftarrow$ mean(kScoresAll);  $\triangleright$ Average Key scores
@@
vScoresMean $\leftarrow$ mean(vScoresAll);  $\triangleright$ Average Value scores
return kScoresMean, vScoresMean;
Appendix C Additional Performance Evaluation of KVmix with LongBench
@@
We present extended experimental results building upon the experiments detailed in Section 4.3.1, aiming to provide a more comprehensive performance comparison and to illustrate the impact of different KVmix configurations on model accuracy. Fig. 8 displays the detailed configuration results when the proportion of high-bit quantization layers is increased to 30%, corresponding to KVmix-k2.28v2.56. Table 4 summarizes the performance comparison of different models using KVmix across various quantization configurations. The experimental results demonstrate that KVmix-k2.19v2.38 achieves a significant accuracy improvement over KVmix-2bit and random-k2.19v2.38. Compared to KVmix-4bit, KVmix-k2.19v2.38 has an average accuracy loss of within 0.013, but achieves nearly twice the KV Cache compression. Further increasing the proportion of high-bit quantization layers to 30% (KVmix-k2.28v2.56) leads to a noticeable improvement in model accuracy, almost the same level as KVmix-4bit, validating the effectiveness and flexibility of KVmix.
@@
Table 4: Model accuracy of 4 LLMs on LongBench with different quantization configurations. KVmix-k2.19v2.38 uses the configurations of Section 4.2. KVmix-k2.28v2.56 uses the configurations of Fig. 8. KVmix-2bit uses the asymmetric 2-bit (Key per-channel and Value per-token) quantization for all model layers (RPC ratio is set to 10%). KVmix-4bit uses the asymmetric 4-bit quantization for all model layers (RPC ratio is set to 20%). random-k2.19v2.38 randomly selects 20% of the model layers to perform asymmetric 3-bit and 4-bit quantization for Key and Value (RPC ratio is set to 20%), and the remaining layers are 2-bit quantization (RPC ratio is set to 10%).

### v2

@@ Conclusion
This paper proposes KVmix, a novel mixed quantization method tackling the KV Cache memory bottleneck in LLM inference. KVmix creatively integrates layer importance analysis based on KV weight gradients into KV quantization and integrates dynamic long-context optimization to cut memory usage while maintaining generation quality. It achieves significant memory and efficiency gains with minimal loss in accuracy, offering flexibility to adapt quantization strategies to diverse scenarios. Future work will explore integrating lightweight mechanisms for real-time KV bit adjustments into KVmix to enhance adaptability.
Acknowledgments
This research was funded by National Key R&D Program of China (2022YFB4501604).
Appendix A
KV Weight Analysis
Fig. 9 provides a comprehensive heatmap analysis of the Key ($W_{k}$) and Value ($W_{v}$) projection weight matrices across different layers for three models: Llama2-7B-hf, Llama 3-8B-Instruct, and Mistral-7B-Instruct-v0.3. Through a multi-model comparison, Fig. 1 reveals the detailed characteristics of the weight distributions for $W_{k}$ and $W_{v}$. The heatmaps highlight the following observations: there are significant differences in the magnitude and distribution of $W_{k}$ and $W_{v}$ across layers in different models, and within the same layer, the weight patterns vary distinctly across models. These findings further validate two key conclusions: (1) the variability of KV weights across layers suggests the need for layer-specific quantization strategies rather than a uniform low-bit quantization strategy; (2) the difference in Key and Value weights distribution within the same layer indicates the need to use different quantization bits for Key and Value. This analysis provides a foundation for the layer importance-aware quantization method, demonstrating the necessity of customizing quantization bit-widths for different layers and models.
@@ Acknowledgments
Figure 9: Projection matrix weights of K and V across different model layers for Llama 2-7B-hf (top), Llama 3-8B-Instruct (middle) and Mistral-7B-Instruct-v0.3 (bottom). The heatmap visualizes the projection weights by sampling every sixth element in the matrices. Additionally, the weight values along the axes have been normalized. “Norm" represents the L2 norm of the weight matrix for each layer, while “Range" indicates the range of values within each layer’s weight matrix.
Appendix B
KVmix Profiler Algorithm
In Algorithm KVmix Profiler Algorithm, we present the algorithm workflow of the KVmix profiler, which encompasses three core steps: data processing, KV importance analysis, and quantization parameters configuration. This process is performed offline. By computing the gradient norms of the model’s loss function with respect to $W_{k}$ and $W_{v}$, the importance of KV at each layer is assessed. Once the layer importance analysis and quantization parameters configuration are completed, the quantization results can be directly reused for subsequent inference tasks without the need for recalculation, thereby ensuring that inference efficiency remains unaffected. Although the analysis involves backpropagation, which demands significant memory, this pressure can be effectively mitigated by limiting the length of input prompts. Selecting a sufficient number of prompts (e.g., 20 to 30) ensures that this optimization does not significantly impact the accuracy of the layer importance analysis.
@@
1:Model modelmodel, inputs input_idsinput\_ids, attention_maskattention\_mask, labelslabels, layers num_layersnum\_layers, datasets datasetsdatasets
2:Quantization bits quant_bitsquant\_bits
3:for dataset in datasets do
4:  prompts $\leftarrow$ randomSample(dataset, n);  $\triangleright$ Sample n prompts
5:  inputs $\leftarrow$ tokenize(prompts);  $\triangleright$ Prepare inputs
6:  model $\leftarrow$ loadFullModel();  $\triangleright$ Load full-precision model
@@
7:  kScores, vScores $\leftarrow$ calcImportance(model, inputs, num_layers);  $\triangleright$ Compute importance
@@
8:  kLayers, vLayers $\leftarrow$ classify(kScores, vScores, num_layers);  $\triangleright$ Classify layers
@@
9:  quantBits $\leftarrow$ setBits(kLayers, vLayers);  $\triangleright$ Set quantization bits
10:  del model;  $\triangleright$ Free memory
11:  model $\leftarrow$ loadQuantModel(quantBits);  $\triangleright$ Load quantized model
12:  preds $\leftarrow$ infer(model, dataset);  $\triangleright$ Inference
13:  save(preds);  $\triangleright$ Save results
14:  del model;  $\triangleright$ Free memory
15:end for
@@
16:Function calcImportance(model, inputs, num_layers):  $\triangleright$ Compute importance scores
17:kScoresAll, vScoresAll $\leftarrow$ [], [];
18:for input in inputs do
19:  kScores, vScores $\leftarrow$ [], [];
20:  for layer in range(num_layers) do
21:   loss $\leftarrow$ model.forward(input);  $\triangleright$ Compute loss
@@
22:   kGrad $\leftarrow$ grad(loss, kParams(layer));  $\triangleright$ Gradient for Key
@@
23:   vGrad $\leftarrow$ grad(loss, vParams(layer));  $\triangleright$ Gradient for Value
24:   kScores.append(kGrad.norm());  $\triangleright$ Append Key score
25:   vScores.append(vGrad.norm());  $\triangleright$ Append Value score
26:   clearMemory();  $\triangleright$ Free memory
27:  end for
28:  kScoresAll.append(kScores);
29:  vScoresAll.append(vScores);
30:end for
31:kScoresMean $\leftarrow$ mean(kScoresAll);  $\triangleright$ Average Key scores
@@
32:vScoresMean $\leftarrow$ mean(vScoresAll);  $\triangleright$ Average Value scores
33:return kScoresMean, vScoresMean;
Appendix C
Analysis Results of KVmix Profiler Using Different Prompts
To verify the robustness of the KVmix profiler, we evaluated its layer importance analysis using various sets of prompts. These included the first 20 and 30 prompts from the LongBench TriviaQA dataset, 20 and 30 prompts randomly selected from LongBench, and 20 and 30 prompts randomly chosen from the Wikitext-2 dataset. The results, as shown in Fig. 10, demonstrate that the distribution of layer importance remains highly consistent regardless of the source or number of prompts used. This consistency is primarily due to our reliance on the model’s layer KV weights ($W_{k}$ and $W_{v}$) as the basis for analyzing KV importance; since the $W_{k}$ and $W_{v}$ remain unchanged during inference, a sufficient number of prompts is adequate to yield reliable importance analysis results. Furthermore, the analysis using 20 prompts yields results nearly identical to those obtained with 30 prompts, with the differences having a negligible impact on the model’s final inference accuracy. This stability confirms the reliability of the KVmix profiler across different input scenarios.
Figure 10: KV configuration results from the KVmix profiler using different datasets and prompts (20% of the layers quantized to 3 bits or 4 bits, and the remaining layers quantized to 2 bits).
Appendix D
Impact of RPC Proportion on Model Accuracy
We systematically evaluated the impact of the Recent Pivotal Context (RPC) proportion on model accuracy by adjusting its proportion in the KVmix-k2.19v2.38 configuration, with results presented in Table 4. The experiments demonstrate that when the RPC ratio is set to 0% (without using RPC), the model’s accuracy decreases by 2.67% compared to the FP16 baseline. In contrast, when employing a 20%/10% RPC ratio, the accuracy only drops by 0.86% relative to the baseline, highlighting the importance of RPC in preserving model accuracy. However, when the RPC ratio exceeds 20%, further increases in the ratio result in only marginal improvements in model accuracy. Additionally, in short-context scenarios, where fewer new KV pairs are generated during the decoding phase, an excessively high RPC ratio results in a large proportion of full-precision KVs, thereby diminishing the memory compression benefits of quantization. As illustrated in Fig. 11, we evaluated the impact of varying RPC ratios on model accuracy and memory compression ratio using the KVmix-k2.19v2.38 configuration on the GSM8K dataset. The results indicate that when the RPC ratio for critical layers exceeds 20%, the model accuracy remains nearly constant, while the memory compression ratio experiences a substantial decline. Therefore, it is recommended to limit the RPC ratio to within 20% when configuring the KVmix profiler to achieve a balance between maintaining inference accuracy and avoiding excessive memory pressure.
Methods Datasets Average
TriviaQA
Qasper
MF-en
QMSum
2WikiMQA
GovRep
Rbench-P
LCC
TREC
PsgRetr-en
FP16 78.89 9.55 22.86 21.19 9.94 17.36 55.64 66.70 66.00 6.64 35.477
w/oRPC 77.95 9.19 21.03 19.98 9.05 15.10 56.13 66.00 65.50 5.61 34.554
10%/0% 78.20 9.29 21.55 19.90 9.33 15.10 56.44 66.30 66.00 5.60 34.771
10%/10% 78.20 9.35 21.53 20.15 9.35 15.10 56.40 66.35 66.00 5.65 34.808
20%/10% 78.78 9.59 22.82 20.49 9.77 15.45 56.54 66.59 66.00 5.72 35.175
20%/20% 78.78 9.63 22.80 20.55 9.77 15.43 56.55 66.50 66.00 5.78 35.179
30%/30% 78.02 10.23 23.21 20.63 9.75 15.60 56.61 66.70 66.00 5.98 35.273
Table 4: Model accuracy of Llama 2-7B on LongBench under different RPC ratios. “w/o RPC": Indicates that the RPC ratio is set to 0 for all layers. “10%/0%": Indicates that the RPC ratio is set to 10% for Keys and Values quantized at 3-bit and 4-bit, and 0% for Keys and Values quantized at 2-bit, with similar interpretations for other ratios.
Figure 11: Accuracy and memory compression ratio variations of KVmix-k2.19v2.38 on Llama 2-7B with GSM8K dataset using different RPC ratios.
Appendix E
Additional Performance Evaluation of KVmix on LongBench
@@
In this section, we present extended experimental results on LongBench to provide a more comprehensive performance comparison and to elucidate the impact of different KVmix configurations on model accuracy. Fig. 12 illustrates the detailed configuration when the proportion of high-bit quantization layers is increased to 30%, corresponding to KVmix-k2.28v2.56. Table 5 supplements the performance data of various models under different KVmix quantization schemes. The results indicate that KVmix-k2.19v2.38 achieves a significant improvement in accuracy compared to KVmix-2bit and random-k2.19v2.38, highlighting the advantages of importance-aware quantization. When compared to KVmix-4bit, KVmix-k2.19v2.38 maintains an average accuracy loss within 1.30% while achieving nearly twice the KV Cache compression ratio. Furthermore, increasing the proportion of high-bit quantization layers to 30% (KVmix-k2.28v2.56) leads to an enhancement in model accuracy, approaching that of KVmix-4bit. These findings validate the flexibility of the KVmix framework in balancing accuracy and efficiency.
@@
Table 5: Model accuracy of 4 LLMs on LongBench with different quantization configurations. KVmix-k2.28v2.56 uses the configurations of Fig. 12. KVmix-2bit uses the asymmetric 2-bit (Key per-channel and Value per-token) quantization for all model layers (RPC ratio is set to 10%). KVmix-4bit uses the asymmetric 4-bit quantization for all model layers (RPC ratio is set to 20%). random-k2.19v2.38 randomly selects 20% of the model layers to perform asymmetric 3-bit and 4-bit quantization for Key and Value (RPC ratio is set to 20%), and the remaining layers are 2-bit quantization (RPC ratio is set to 10%).